# Magnetic Localization for In-body Nano-communication Medical Systems

### v1

@@ Abstract
Nano-machines circulating inside the human body, collecting data on tissue conditions, represent a vital part of next-generation medical diagnostic systems. However, for these devices to operate effectively, they need to relay not only their medical measurements but also their positions. This paper introduces a novel localization method for in-body nano-machines based on the magnetic field, leveraging the advantageous magnetic permeability of all human tissues. The entire proposed localization system is described, starting from 1010 magnetometers to be integrated into the nano-machines, to a set of external wires generating the magnetic field. Mathematical equations for the localization algorithm are also provided, assuming the nano-machines do not execute the computations themselves, but transmit their magnetic field measurements together with medical data outside of the body. The whole system is validated with computer simulations that capture the measurement error of the magnetometers, the error induced by the Earth’s magnetic field, and a human body model assuming different possible positions of nano-machines. The results show a very high system accuracy with localization errors even below 1 cm.
@@
Flow-guided nano-networks, IoT for health, magnetic field, nano-communications, THz communications, wireless localization.I Introduction
@@
Having this in mind, the contribution of this paper is to propose a novel localization approach for in-body nano-machines based on the magnetic field. Like in many localization systems with reference (anchor) points, the reference wires are generating a constant magnetic field. These wires are located in known positions outside of the human body and the emitted magnetic field is constant in time but decreases with the distance from the wires. As the magnetic field penetrates human tissues very well (the relative magnetic permeability is very close to 1 for all human tissues), it can be exploited for determining the distances of a nano-machine from the wires and later the nano-machine position. Moreover, in the proposed solution, the nano-machines do not perform any computations. They just measure the magnetic field, along with other medical measurements. Later, they transmit both the medical and magnetic field measurements through some intermediate devices to a system outside of the body where all the calculations are performed. We present the whole localization system, starting from the 1010 magnetometers to be integrated into the nano-machines, to the set of wires generating the magnetic field. We provide the equations for the localization algorithm in two versions: for three wires and nine (or more) wires. Then, we assess the localization accuracy with computer simulations taking into account the measurement error of the magnetometers, the error induced by the Earth’s magnetic field, and a human body model assuming different possible positions of nano-machines. The results show a very good system accuracy with localization error even below 1 cm.
@@
II Related Work on In-body Nano-node Localization
@@
Performing wireless localization of mobile nano-nodes inside a human body poses several additional challenges compared with the localization in macro-scale wireless networks. First, radio propagation conditions inside the body are quite harsh, with very high signal attenuation, strongly dependent on the type of tissue. For example, for the frequency of 1 THz, which is typically considered for EM nano-nodes, the signal propagation loss at a distance of 1 mm is about 70 dB for fat tissue, 90 dB for skin and even 150 dB for blood [19], which makes classical receive signal strength (RSS) localization approaches hardly applicable. Second, this extremely high signal attenuation limits the communication range of nano-nodes to about 1-2 mm, so a nano-node flowing with blood is frequently isolated from other nodes, unless it approaches another nano-device.
@@
The number of already proposed localization schemes for nano-networks is very limited. Considering nano-nodes flowing with blood in a cardiovascular system, an inertial positioning system has been recently proposed [20]. In this paper, nano-nodes are assumed to get information about their estimated position from anchor nodes attached to the skin, when they flow close by. Later, during their movement, nano-nodes update their positions based on their velocity and rotation sensors. However, this concept is very challenging, as the blood flow is not laminar all the time, but also turbulent, which makes the position updates inaccurate. Moreover, communication between the skin-attached anchors and nano-nodes is rather not possible directly, because of the distance range usually about a centimeter or more. In another paper [21], the time of flight approach was suggested to calculate the two-way distance between anchors and nano-nodes and later to perform localization. This was, though, considered for a grid of static nano-nodes. This concept was later developed in [22], also for static grid-type scenario, where nano-nodes further away from the anchors were localized based on the positions of nano-nodes closer to the anchors. Two other research contributions [23, 24] focus on the hop-counting approach, but this is again feasible only for static nano-nodes arranged in a grid.
@@
Having in mind these discussed limitations of wireless localization, here, we propose to exploit the properties of a static magnetic field. Its main advantage is the invariability in different types of human tissues. The magnetic permeability, which describes how the static magnetic field penetrates different media, is practically the same, up to the sixth decimal place, in all human tissues, see Tab. I. This means that magnetic field strength values can be used for distance estimation independently on the tissues between a magnetic field transmitter and a receiver, e.g., a nano-node.
@@
To the best of the authors’ knowledge, this is the first paper where magnetic field localization is proposed for nano-nodes. There are, however, some published works discussing localization with magnetic fields for networks of larger scale. Localization with the magnetic field generated by external magnets is discussed in [26]. The proposed system might be applied inside a body, but the designed mm-size chips transmitting at 480-520 MHz are too large for nano-networks. A system of magnetic coils working as anchors is considered in [27, 28, 29]. Also there, the in-body devices are not suitable for nano-networks, as both the magnetic sensor and the wireless transponder are mm-size. Magnetic field localization is proposed as well for applications with endoscopy capsules being cm-size, assuming a solenoid [30] or a soft-magnet [31] generating the field from inside of a body and some magnetic sensors on the body. In-body magnet rotation can be also tracked for such endoscopy scenarios [32]. Finally, a thorough review of indoor and outdoor magnetic field positioning systems is given in [33].
III Magnetic Localization System
Having in mind the main considered application, i.e., in-body nano-networks, here we propose a magnetic field localization system composed of (a) reference anchors outside of a body, and (b) tiny magnetometers mounted on boards of in-body nano-machines. We also assume that information from nano-nodes is transmitted outside of the body through some intermediate devices like a nano-gateway. The whole localization process starts with reference anchors emitting the magnetic field. Then, the nano-nodes measure the magnetic field with their magnetometers. The nano-nodes do not perform any computations. Instead, they transmit these readings, together with other medical measurements, through the intermediate devices to the outside of the body. This approach keeps the nano-nodes simple and avoids wasting their precious energy for localization data processing. The concept is presented in Fig. 1 and the main system components are discussed in the three following subsections.
@@ III-A Reference Anchors
In localization systems, anchors act as sources of reference signals that enable a localized device to calculate the distances from the anchors. In a magnetic field localization system, the anchors should generate a magnetic field. It has been already suggested to use some coils for this purpose, working like electromagnets [28]. However, the mathematical equations of that solution do not have a closed form and they have to be solved numerically with some initial guess values. Instead, here we propose to use a set of metallic wires, each of them supplied with a constant electric current (direct current DC). According to the Biot-Savart law, such a DC wire generates a constant magnetic field decreasing with the distance from the wire axis. Thus, by measuring the magnetic field, a distance from the wire can be determined. It is the basic principle of the localization algorithm proposed here; its mathematical equations are given in Sec. IV. We present two versions of the algorithm, the first one for just three wires, and the second one, more robust and accurate, for nine or more wires.
@@ III-B Nano-machines with Magnetometers
Mobile nano-nodes form a nano-network operating inside the body, e.g., in the cardiovascular system, collecting medical information relating to different pathologies, from tissue damage to cancer biomarkers [35]. Each nano-node, of size about 10 , is a very simple embedded system powered with its own energy source (e.g., a piezoelectric generator) and limited processing, data storage, and communication capabilities [7, 36].
@@
In this paper, we assume the nano-nodes, in addition to medical sensors, are equipped with magnetometers, i.e., instruments to measure the magnetic field. In general, magnetometers are large-scale devices that can be classified into many types like magnetoresistive, spin-valve, superconducting quantum interference devices, or Hall effect-based. For nano-nodes, the magnetometer size should be much smaller than most of them, so here we propose to take advantage of a very recently proposed solution: a graphene Hall effect magnetometer [37]. The active area of this magnetometer is 10 x 10 , which is comparable with state-of-the-art nano-nodes and hopefully could be even reduced with the progress in this technology. This type of device was already tested experimentally, reporting a steady relative measurement error of 1% for the magnetic field up to 120 mT [37].
@@
The magnetic field is a vector parameter and a single Hall effect magnetometer measures just a single component: , , or . Thus, each nano-node is equipped with three such devices, together measuring all three field components.
@@
When performing a medical measurement that requires additional information where the measurement is taken, the magnetometers’ readings are saved together with the medical ones. Later, all this data is transmitted out of the body through the intermediate devices. The nano-node position is calculated after that, in a medical data center or a doctor’s computer, so that the nano-node processor is not burdened with these calculations.
@@ III-C Intermediate Devices
As mentioned in Sec. II, the tiny nano-nodes circulating in a human body have very restricted capabilities, with a communication range limited to about 1-2 mm. Because of that, a nano-node flowing through the same vein where a nano-gateway is installed has only a limited probability of successful transmission to this device [38]. This topic was already further investigated for so-called flow-guided nano-networks in numerous medical applications like diagnosis of artery occlusion [39], bacterial infections, sepsis, heart attacks, or restenosis [40]. The results show that with a sufficient number of nano-nodes in the network, the probability of a successful information transfer out of the body is close to 100%, but it may take even a few hours [39]. A recent work [41] shows that multi-hop communication between the nano-nodes may additionally improve this information transfer. While the detailed analysis of this communication is out of the scope of this paper, we assume that nano-nodes can deliver their gathered medical and magnetic field data to a medical data center through intermediate devices like a nano-router and a body area network.
@@ IV Localization Algorithm
In this section, we explain the proposed localization algorithm. Our proposed solution shares some similarities with the receive signal strength approach, but here the magnetic flux density is measured instead of radio signal strength. The proposed algorithm has two phases: ranging and lateration. While the ranging is the same for different arrangements of DC wires, the lateration phase depends on the number of wires. The basic approach requires only three wires, one for each plane. The more complex algorithm requires at least nine wires, three for each plane.
@@ IV-A Ranging
In the ranging phase, each DC wire is activated in a sequence. At any moment, only one wire is activated. First, nano-sensors measure the magnetic flux density generated by all DC wires, one wire at a time. Since each nano-sensor has three magnetometers, three measurements are taken simultaneously for each wire, corresponding to three orthogonal components of the flux density. These three measurements are then combined into a single vector, whose magnitude is calculated. Finally, the distance from the wire is calculated with the Biot-Savart law:
@@ IV-B Trilateration with Minimal Number of Wires
In this scenario, the system is considered to have the minimal required number, i.e., three wires. The wires are aligned along the X, Y, and Z axes of the coordinated system (see the wires X, Y, and Z0 in Fig. 2). The nano-nodes are located in the (+x, +y, +z) octant of the coordinate system, so all three coordinates are always positive values. Each distance between a nano-node and a wire, calculated in the ranging phase, creates a cylinder of possible nano-node positions along the wire. For each wire, only two coordinates are involved. Thus, with three wires, a set of simple three equations can be formed, one for each plane:
@@
, , - unknown coordinates of the nano-node
, , - distances from the nano-node to the wires X, Y, Z, respectively.
@@
While this approach theoretically can provide the nano-node position, in practice, it is not robust against measurement errors. When a nano-node is located close to one of the , , or planes (see Fig. 2), one of the estimated distances might be much larger than the other ones. Consequently, one of the root values in (3) might be negative. It is then set to zero, to avoid false imaginary parts of the calculated coordinates.

### v2

@@ Abstract
Nano-machines circulating inside the human body, collecting data on tissue conditions, represent a vital part of next-generation medical diagnostic systems. However, for these devices to operate effectively, they need to relay not only their medical measurements but also their positions. This paper introduces a novel localization method for in-body nano-machines based on the magnetic field, leveraging the advantageous magnetic permeability of all human tissues. The entire proposed localization system is described, starting from 10 10 magnetometers to be integrated into the nano-machines, to a set of external wires generating the magnetic field. Mathematical equations for the localization algorithm are also provided, assuming the nano-machines do not execute the computations themselves, but transmit their magnetic field measurements together with medical data outside of the body. The whole system is validated with computer simulations that capture the measurement error of the magnetometers, the error induced by the Earth’s magnetic field, and a human body model assuming different possible positions of nano-machines. The results show a very high system accuracy with position errors even below 1 cm.
@@
Flow-guided nano-networks, IoT for health, magnetic field, nano-communications, sensors, THz communications, wireless localization.I Introduction
@@
Having this in mind, the contribution of this paper is to propose a novel localization approach for in-body nano-machines based on the magnetic field. Like in many localization systems with reference (anchor) points, the reference wires are generating a constant magnetic field. These wires are located in known positions outside of the human body and the emitted magnetic field is constant in time but decreases with the distance from the wires. As the magnetic field penetrates human tissues very well (the relative magnetic permeability is very close to 1 for all human tissues), it can be exploited for determining the distances of a nano-machine from the wires and later the nano-machine position. Moreover, in the proposed solution, the nano-machines do not perform any computations. They just measure the magnetic field, along with other medical measurements. Later, they transmit both the medical and magnetic field measurements wirelessly, e.g., using the THz band [9], through some intermediate devices to a system outside of the body where all the calculations are performed. We present the whole localization system, starting from the 10 10 magnetometers to be integrated into the nano-machines, to the set of wires generating the magnetic field. We provide the equations for the localization algorithm in two versions: for three wires and six or more wires. Then, we assess the localization accuracy with computer simulations taking into account the measurement error of the magnetometers, the error induced by the Earth’s magnetic field, and a human body model assuming different possible positions of nano-machines. The results show a very good system accuracy with localization error even below 1 cm.
@@
II Related Work on In-body Nano-machine Localization
@@
Performing wireless localization of mobile nano-machines inside a human body poses several additional challenges compared with the localization in macro-scale wireless networks. First, radio propagation conditions inside the body are quite harsh, with very high signal attenuation, strongly dependent on the type of tissue. For example, for the frequency of 1 THz, which is typically considered for EM nano-nodes, the signal propagation loss at a distance of 1 mm is about 70 dB for fat tissue, 90 dB for skin and even 150 dB for blood [19], which makes classical receive signal strength (RSS) localization approaches hardly applicable. Second, this extremely high signal attenuation limits the communication range of nano-nodes to about 1-2 mm, so a nano-machine flowing with blood is frequently isolated from other nodes, unless it approaches another nano-machine.
@@
The number of already proposed localization schemes for nano-networks is very limited. Considering nano-machines flowing with blood in a cardiovascular system, an inertial positioning system has been recently proposed [20]. In this paper, nano-machines are assumed to get information about their estimated position from anchor nodes attached to the skin, when they flow close by. Later, during their movement, nano-machines update their positions based on their velocity and rotation sensors. However, this concept is very challenging, as the blood flow is not laminar all the time, but also turbulent, which makes the position updates inaccurate. Moreover, communication between the skin-attached anchors and nano-machines is rather not possible directly, because of the distance range usually about a centimeter or more. In another paper [21], the time of flight approach was suggested to calculate the two-way distance between anchors and nano-machines and later to perform localization. This was, though, considered for a grid of static nano-machines. This concept was later developed in [22], also for static grid-type scenario, where nano-machines further away from the anchors were localized based on the positions of nano-machines closer to the anchors. Two other research contributions [23, 24] focus on the hop-counting approach, but this is again feasible only for static nano-machines arranged in a grid.
@@
Having in mind these discussed limitations of wireless localization, here, we propose to exploit the properties of a static magnetic field. While an alternating magnetic field, which is associated with electromagnetic waves propagation, is strongly influenced by the type of human tissues it passes through, the situation is different with a static magnetic field. The main advantage of the static magnetic field is its invariability in different types of human tissues. The magnetic permeability, which describes how the static magnetic field penetrates different media, is practically the same, up to the sixth decimal place, in all human tissues [25, 26, 27, 28], see Tab. LABEL:TablePermeability. This means that magnetic field strength values can be used for distance estimation independently on the tissues between a magnetic field transmitter and a receiver, e.g., a nano-machine. Moreover, since the relative permeability of various body tissues is close to 1, the static magnetic field experiences minimal disturbance as it moves through a body. It is an additional advantage which facilitates any kind of in-body magnetic field measurements.
To the best of the authors’ knowledge, this is the first paper where magnetic field localization is proposed for nano-machines. There are, however, some published works discussing localization with magnetic fields for networks of larger scale. Localization with the magnetic field generated by external magnets is discussed in [29]. The proposed system might be applied inside a body, but the designed mm-size chips transmitting at 480-520 MHz are too large for nano-networks. A system of magnetic coils working as anchors is considered in [30, 31, 32]. Also there, the in-body devices are not suitable for nano-networks, as both the magnetic sensor and the wireless transponder are mm-size. Magnetic field localization is proposed as well for applications with endoscopy capsules being cm-size, assuming a solenoid [33] or a soft-magnet [34] generating the field from inside of a body and some magnetic sensors on the body. In-body magnet rotation can be also tracked for such endoscopy scenarios [35]. Finally, a thorough review of indoor and outdoor magnetic field positioning systems is given in [36].
III Magnetic Localization System
Having in mind the main considered application, i.e., in-body nano-networks, here we propose a magnetic field localization system composed of (a) reference anchors outside of a body, and (b) tiny magnetometers mounted on boards of in-body nano-machines. We also assume that information from nano-machines is transmitted outside of the body through some intermediate devices like a nano-gateway. The whole localization process starts with reference anchors emitting the magnetic field. Then, the nano-machines measure the magnetic field with their magnetometers. The nano-machines do not perform any computations. Instead, they wirelessly transmit these readings, together with other medical measurements, through the intermediate devices to the outside of the body. There, the readings are used to calculate the distances of each nano-machine from the wires. Finally, the 3D position of the nano-machine is determined with the aid of the localization algorithm (trilateration or multilateration). This approach keeps the nano-machines simple and avoids wasting their precious energy for localization data processing. The concept is presented in Fig. 1 and the main system components are discussed in the three following subsections.
@@ III-A Reference Anchors
In localization systems, anchors act as sources of reference signals that enable a localized device to calculate the distances from the anchors. In a magnetic field localization system, the anchors should generate a magnetic field. It has been already suggested to use some coils for this purpose, working like electromagnets [31]. However, the mathematical equations of that solution do not have a closed form and they have to be solved numerically with some initial guess values. Instead, here we propose to use a set of metallic wires, each of them supplied with a constant electric current (direct current DC). According to the Biot-Savart law, such a DC wire generates a constant magnetic field decreasing with the distance from the wire axis. Thus, by measuring the magnetic field, a distance from the wire can be determined. It is the basic principle of the localization algorithm proposed here; its mathematical equations are given in Sec. IV. We present two versions of the algorithm, the first one for just three wires, and the second one, more robust and accurate, for six or more wires.
@@ III-B Nano-machines with Magnetometers
Mobile nano-machines form a nano-network operating inside the body, e.g., in the cardiovascular system, collecting medical information relating to different pathologies, from tissue damage to cancer biomarkers [38]. Each nano-machine, of size about 10 3, is a very simple embedded system powered with its own energy source (e.g., a piezoelectric generator) and limited processing, data storage, and communication capabilities [7, 39].
The miniaturization of an antenna to meet the size requirements of a nano-machine imposes the utilization of very high resonant frequencies. For example, a one-micrometer-long antenna built with a perfect electric conductor (PEC) material would resonate at 150 THz, i.e., in the infrared optical region of the electromagnetic spectrum. There are several challenges associated with the development of optical antennas, starting from the fact that there are no real PEC materials and even very good conductors such as gold or silver have finite complex-valued conductivity, which impacts the antenna design and performance [40, 41]. When utilizing optical antennas, on-chip lasers and nano-photodetectors are needed [42, 43].
Alternatively, to operate at lower frequencies, graphene-based plasmonic transceivers and antennas could be utilized. Indeed, graphene supports the propagation of surface plasmon polariton waves at terahertz-band frequencies (i.e., between 100 GHz and 10 THz), with very high confinement factors. As a result, a one-micrometer-long graphene-based antenna radiates at approximately 2 THz, i.e., at a frequency nearly two orders of magnitude lower than that of a metallic antenna [44, 45]. When utilizing such antennas, a graphene-based plasmonic nano-transceiver is needed to generate in transmission and process in reception THz signals [46, 47]. If even lower frequencies are needed, recently, magneto-electric antennas able to operate at hundreds of MHz have been proposed [48, 49]. In all these cases, individual elements have been developed, and their integration in a miniature nano-radio is the next step. In [50], a good summary of the state of the art is provided.
When it comes to the energy management system of nano-machines, the situation is quite similar. Due to the nature of nano-machines and the inability to manually replace or recharge their batteries, energy harvesting systems are critical [51]. Different technologies have been proposed to collect energy from the environment. The most commonly cited relies on piezoelectric nanogenerators built with zinc oxide nanowires that harvest energy from the heartbeat or the blood flow [52]. There have been several studies aimed at optimizing the performance of self-powered nano-machines [53].
@@
In this paper, we assume the nano-machines, in addition to medical sensors, are equipped with magnetometers, i.e., instruments to measure the magnetic field. In general, magnetometers are large-scale devices that can be classified into many types like magnetoresistive, spin-valve, superconducting quantum interference devices, or Hall effect-based. For nano-machines, the magnetometer size should be much smaller than most of them, so here we propose to take advantage of a very recently proposed solution: a graphene Hall effect magnetometer [54]. The active area of this magnetometer is 10 x 10 , which is comparable with state-of-the-art nano-machines and hopefully could be even reduced with the progress in this technology. This type of device was already tested experimentally, reporting a steady relative measurement error of 1% for the magnetic field from 0 to 120 mT, with the step of 2 mT [54].
@@
The magnetic field is a vector parameter and a single Hall effect magnetometer measures just a single component: , , or . Thus, each nano-machine is equipped with three such devices, together measuring all three field components.
In addition to the magnetometers, a control unit, a communication unit and an energy management system, at the very least, are needed. Aligned with the vision of nano-machines [7, 39, 50], the size of the envisioned embedded nano-systems should not exceed tens to a few hundred cubic micrometers. The energy management nano-system, often considered to be based on piezoelectric or acoustic harvesting technologies, has usually been the dominating element. Instead, compact on-chip radios based either on plasmonic or magnetoelectric antennas are significantly smaller (sub-micrometric). With current experimentally developed solutions, the magnetometers will be comparable in size to the energy harvesting nano-system. This motivates further research on nano-systems integration.
@@
When performing a medical measurement that requires additional information where the measurement is taken, the magnetometers’ readings are saved together with the medical ones. Later, all this data is transmitted out of the body through the intermediate devices. The nano-machine position is calculated after that, in a medical data center or a doctor’s computer, so that the nano-machine processor is not burdened with these calculations.
@@ III-C Intermediate Devices
As mentioned in Sec. II, the tiny nano-machines circulating in a human body have very restricted capabilities, with a communication range limited to about 1-2 mm. Because of that, a nano-machine flowing through the same vein where a nano-gateway is installed has only a limited probability of successful transmission to this device [55]. This topic was already further investigated for so-called flow-guided nano-networks in numerous medical applications like diagnosis of artery occlusion [56], bacterial infections, sepsis, heart attacks, or restenosis [57]. The results show that with a sufficient number of nano-machines in the network, the probability of a successful information transfer out of the body is close to 100%, but it may take even a few hours [56]. A recent work [58] shows that multi-hop communication between the nano-machines may additionally improve this information transfer. While the detailed analysis of this communication is out of the scope of this paper, we assume that nano-machines can deliver their gathered medical and magnetic field data to a medical data center through intermediate devices like a nano-router and a body area network. This communication process is assumed to be performed after the localization procedure, which takes no longer than 1 ms (see Sec. V-B).
@@ IV Localization Algorithm
In this section, we explain the proposed localization algorithm. Our proposed solution shares some similarities with the receive signal strength approach, but here the magnetic flux density is measured instead of the radio signal strength. The proposed algorithm has two phases: ranging and lateration. While the ranging is the same for different arrangements of DC wires, the lateration phase depends on the number of wires. The basic approach requires only three wires, one for each plane. The more complex algorithm requires at least six wires, three in one plane and three in another plane.
@@ IV-A Ranging
In the ranging phase, each DC wire is activated in a sequence. At any moment, only one wire is activated. First, nano-machines measure the magnetic flux density generated by all DC wires, one wire at a time. Since each nano-machine has three magnetometers, three measurements are taken simultaneously for each wire, corresponding to three orthogonal components of the flux density. These three measurements are then combined into a single vector, whose magnitude is calculated. Finally, the distance from the wire is calculated with the Biot-Savart law:
@@ IV-B Trilateration with Minimal Number of Wires
In this scenario, the system is considered to have the minimal required number, i.e., three wires. The wires are aligned along the X, Y, and Z axes of the coordinated system (see the wires X, Y, and Z0 in Fig. 2). The nano-machines are located in the (+x, +y, +z) octant of the coordinate system, so all three coordinates are always positive values. Each distance between a nano-machine and a wire, calculated in the ranging phase, creates a cylinder of possible nano-machine positions along the wire. For each wire, only two coordinates are involved. Thus, with three wires, a set of simple three equations can be formed, one for each plane:
@@
, , - unknown coordinates of the nano-machine
, , - distances from the nano-machine to the wires X, Y, Z, respectively.
@@
While this approach theoretically can provide the nano-machine position, in practice, it is not robust against measurement errors. When a nano-machine is located close to one of the , , or planes (see Fig. 2), one of the estimated distances might be much larger than the other ones. Consequently, one of the root values in (3) might be negative. It is then set to zero, to avoid false imaginary parts of the calculated coordinates.